Working together, computer A and B can complete a data-processing job in 2 hours. Computer A working alone can do the job in 3 hours less that computer B working alone. How long does it take each computer to do the job by itself?
Answer by stanbon(75887)

Working together, computer A and B can complete a data-processing job in 2 hours. Computer A working alone can do the job in 3 hours less than computer B working alone. How long does it take each computer to do the job by itself?
----------------
Together rate:: 1/2 job/hr
---------
A rate:: 1/(x-3) job/hr
---------
B rate:: 1/x job/hr
--------------------------
Equation:
rate + rate = together rate
1/(x-3) + 1/x = 1/2
-----
2x + 2(x-3) = x(x-3)
2x + 2x-6 = x^2 - 3x
x^2 - 7x + 6 = 0
---
(x-6)(x-1) = 0
Acceptable answer:: x = 6 hr (B time to do the job)
x-3 = 3 hrs (A time to do the job)
-----------
Cheers,
Stan H.
-----------